eviant:

The domain for f(x) and g(x) is all real numbers. Let f(x) = x2 + x-20 and g(x) = 3x-12 Find g(x)/f(x)and state its domain.

eviant:

@AngeI @Shadow

dude:

Oh simply just put each equation on either the numerator and denominator given the problem \(\dfrac{g(x)}{f(x)}-> \dfrac{3x-12}{x^2+x-20}\)

dude:

Now we know that the denominator cannot equal 0 So set the denominator equal to 0 to see which values are not within the range

eviant:

@dude not sure what you mean by setting the denominator equal to 0?

eviant:

@Vocaloid

dude:

\(x^2 + x-20=0\) Solve for x

eviant:

@dude x=4, -5?

dude:

Other way around sign -4 and 5 So you domain would be all real number except -4 and 5

eviant:

@dude what's g(x)/f(x)?

dude:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dude Oh simply just put each equation on either the numerator and denominator given the problem \(\dfrac{g(x)}{f(x)}-> \dfrac{3x-12}{x^2+x-20}\) \(\color{#0cbb34}{\text{End of Quote}}\)

eviant:

@mhchen @Vocaloid @dude Mathway says the answer is 3/x+5, but idk how to get there

mhchen:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dude \(\color{#0cbb34}{\text{Originally Posted by}}\) @dude Oh simply just put each equation on either the numerator and denominator given the problem \(\dfrac{g(x)}{f(x)} = \dfrac{3x-12}{x^2+x-20}\) \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{End of Quote}}\) \[=\frac{3(x-4)}{(x-4)(x+5)}\]

eviant:

@mhchen thank you so much, I got it now